OpenStudy (anonymous):

Solve the following equation: 5x+sqrtx+7 - 13=0

OpenStudy (anonymous):

x^2+5x-6=0 (x+6)(x-1)=0 x=-6,1

OpenStudy (anonymous):

The book states the answer is x=2

OpenStudy (anonymous):

After you get the values of x. You plug them in the equation see which value is true:)